# Class-Independent Increment: An Efficient Approach for Multi-label Class-Incremental Learning

### v1

@@ IV-A Experimental Setup and Implementation Details
Compared Method. Our method was compared to multiple SOTA continual learning methods following the MLCIL protocol. This included the SOTA MLCIL methods KRT [11] and APPLE [12], along with various single-label class-incremental learning methods for comparison. These single-label methods comprised regularization-based approaches such as EWC [36], LwF [38], and rehearsal-based methods like ER [ER], TPCIL [22], DER++[24], and PODNet[23]. Additionally, we considered prompt-based methods L2P [7] and DualPrompt [17], both utilizing pre-trained ViT-B models and representing the latest SOTA methods in the incremental learning field. The Upper-bound joint trains all classes simultaneously. However, it is not an Incremental Learning (IL) method but serves as an upper bound for performance comparison.
@@
Evaluation metrics. Following the precedent set by prior incremental learning studies [47, 11, 26], we report two key metrics, average accuracy and final accuracy. For a comprehensive evaluation of all learned categories within each session, we utilize the mean average precision (mAP) and report both the average mAP (the mean of the mAP across all sessions) and the final mAP (mAP in the final session). To further enhance the evaluation of performance across all incremental tasks, we include the per-class F1 measure (CF1) and overall F1 measure (OF1) alongside the final accuracy.
@@
Implementation details. All methods were implemented using PyTorch and trained on 2 RTX 3090 GPUs. We employed a TResNetM backbone pre-trained on ImageNet-21k (L2P [7] and Dual-prompt utilized ViT-B/16 pre-trained on ImageNet-21k as the backbone). The input resolution is set to $h\times w=224\times 224$ or $h\times w=448\times 448$ depending on the training protocol. The projection dimension $D$ and hidden dimension $l$ of the CINet module were configured as $768$. Our method was trained for 20 epochs using the Adam optimizer and OneCycleLR scheduler, with a weight decay of $1\times 10^{-4}$. The batch size was set to 64. For training the base model, the learning rate was set to $4\times 10^{-5}$. During the incremental sessions, we set the learning rate to $1\times 10^{-4}$ for COCO and $8\times 10^{-5}$ for VOC. For the hyper-parameters of total optimization, $\alpha$ is typically set to 0.05, $\beta$ is set to 80, and $\lambda$ is 0.1. Data augmentation techniques included rand augmentation and cutout. Additionally, we conducted experiments three times and reported the average results.
TABLE II: Class-incremental results on MS-COCO and PASCAL VOC datasets. Compare methods under different protocols against comparison methods. All the input size is $224\times 224$.
Method Buffer MS-COCO B0-C20 MS-COCO B40-C5 Buffer VOC B0-C4 VOC B10-C2
Size Avg. Acc Last Acc Avg. Acc Last Acc Size Avg. Acc Last Acc Avg. Acc Last Acc
mAP (%) mAP (%) mAP (%) mAP (%) mAP (%) mAP (%) mAP (%) mAP (%)
Upper-bound - - 81.8 - 81.8 - - 93.6 - 93.6
FT [61] 0 51.4 27.0 ($\downarrow 49.7$) 24.5 10.7 ($\downarrow 58.2$) 0 82.4 62.9 ($\downarrow 21.1$) 70.1 43.0 ($\downarrow 31.3$)
KRT [11] 79.4 74.8 ($\downarrow 1.9$) 73.5 65.4 ($\downarrow 3.5$) 89.1 80.2 ($\downarrow 3.8$) 84.3 71.5 ($\downarrow 2.8$)
Ours 80.7 76.7 ($\downarrow 0.0$) 75.4 68.9 ($\downarrow 0.0$) 91.1 84.0 ($\downarrow 0.0$) 85.6 74.3 ($\downarrow 0.0$)
iCaRL [6] 20/class 69.0 52.4 ($\downarrow 24.6$) 62.4 53.9 ($\downarrow 20.6$) 2/class 87.2 72.4 ($\downarrow 14.0$) 79.0 66.7 ($\downarrow 16.8$)
BiC [28] 70.3 51.5 ($\downarrow 25.5$) 60.2 51.7 ($\downarrow 22.8$) 86.8 72.2 ($\downarrow 14.2$) 81.7 69.7 ($\downarrow 13.8$)
ER [25] 68.7 58.0 ($\downarrow 19.0$) 65.5 58.1 ($\downarrow 16.4$) 86.1 71.5 ($\downarrow 14.9$) 81.5 68.6 ($\downarrow 14.9$)
TPCIL [22] 73.2 63.4 ($\downarrow 13.6$) 69.5 64.2 ($\downarrow 10.3$) 87.6 77.3 ($\downarrow 9.1$) 80.7 70.8 ($\downarrow 12.7$)
PODNet [23] 74.9 66.9 ($\downarrow 10.1$) 66.8 58.9 ($\downarrow 15.6$) 88.1 76.6 ($\downarrow 9.8$) 81.2 71.4 ($\downarrow 12.1$)
DER++ [24] 75.9 68.2 ($\downarrow 8.8$) 68.9 62.1 ($\downarrow 12.4$) 87.9 76.1 ($\downarrow 10.3$) 82.3 70.6 ($\downarrow 12.9$)
KRT [11] 79.4 75.2 ($\downarrow 1.8$) 76.3 72.4 ($\downarrow 2.1$) 90.7 83.4 ($\downarrow 3.0$) 87.7 80.5 ($\downarrow 3.0$)
Ours 80.8 77.0 ($\downarrow 0.0$) 78.2 74.5 ($\downarrow 0.0$) 92.4 86.4 ($\downarrow 0.0$) 90.3 83.5 ($\downarrow 0.0$)
TABLE III: Class-Incremental results (mAP%) on MS-COCO dataset against prompt-based CIL method. $\nabla$ indicates the gap towards the Upper Bound of the corresponding backbone. All the input size is $224\times 224$ and all metrics are in %.
@@
L2P [7] 73.28 ($\downarrow 6.23$) 67.72 ( $\nabla$ 15.44 ) 73.07 ($\downarrow 6.66$) 70.42 ( $\nabla$ 12.74 )
L2P-R [7] 73.87 ($\downarrow 5.64$) 68.22 ( $\nabla$ 14.94 ) 73.64 ($\downarrow 6.09$) 71.68 ( $\nabla$ 11.48 )
Dual-prompt [17] 74.67 ($\downarrow 4.84$) 69.39 ( $\nabla$ 13.77 ) 74.45 ($\downarrow 5.28$) 71.95 ( $\nabla$ 11.21 )
Dual-prompt-R [7] 74.87 ($\downarrow 4.64$) 70.20 ( $\nabla$ 12.96 ) 74.54 ($\downarrow 5.19$) 72.60 ( $\nabla$ 10.56 )
Upper-bound TResNetM 29.4M - 81.80 - 81.80
KRT [11] 74.64 ($\downarrow 4.87$) 65.94 ( $\nabla$ 15.86 ) 77.83 ($\downarrow 1.90$) 74.02 ( $\nabla$ 7.78 )
KRT-R [11] 76.53 ($\downarrow 2.98$) 71.24 ( $\nabla$ 10.56 ) 78.34 ($\downarrow 1.39$) 75.18 ( $\nabla$ 6.62 )
Our 78.14 ($\downarrow 1.37$) 71.36 ( $\nabla$ 10.44 ) 78.99 ($\downarrow 0.74$) 75.32 ( $\nabla$ 6.48 )
Our-R 79.51 ($\downarrow 0.00$) 74.47 ( $\nabla$ 7.33 ) 79.73 ($\downarrow 0.00$) 77.14 ( $\nabla$ 4.66 )
IV-B Comparison with State-of-the-arts.
@@
Result on COCO benchmark: Tab I presents the results of the MS-COCO B0-C10 and B40-C10 benchmark. Our method consistently outperforms all comparison methods in terms of average accuracy (mAP) and final accuracy (CF1, OF1, and mAP). Specifically, when the buffer size is larger (20 per class), our method achieves the best final accuracies of 74.5% and 77.1% on the two benchmarks, surpassing SOTA CIL methods by 11.4% and 10.6%, respectively. It also outperforms the latest MLCIL method, KRT, by 4.3% and 1.9%. When the buffer size decreases (5 per class), our shows larger performance gains compared to other continuals. Notably, when the buffer size is set to 0, our maintains superior performance, far exceeding regularization-based methods and other rehearsal-based methods, with improvements of 42.5% and 45.8% on the B0-C10 benchmark, respectively. Furthermore, our method is more suitable for rehearsal-free scenarios. Specifically, under the B0-C10 protocol, our method achieves average and final accuracy of 78.1% and 71.4%, respectively, compared to KRT (buffer size=0), representing improvements of 3.5% and 5.5%, and even outperforming KRT-R (buffer size=20) by 1.6% and 1.2%.
Tab II demonstrates the results for the other two benchmarks of COCO, including B0-C20 and B40-C5. Similar to the previous benchmarks of COCO, our method achieves the best accuracy across multiple metrics. Specifically, it attains final accuracies of 77.0% and 74.5% on the two benchmarks, outperforming rehearsal-based methods by at least 8.8% and 10.3%. Additionally, we observe improvements compared to the KRT method, especially when the buffer size is set to 0. Furthermore, under the relatively simpler protocols, our final accuracy has reached 77.0%, approaching the upper bound of 81.8%, highlighting the effectiveness of our method.
@@
The current SOTA methods in SLCIL are prompt-based approaches. These methods, based on the VIT-B pre-trained model and utilizing prompts, demonstrate powerful capabilities in preventing catastrophic forgetting. For a more comprehensive comparison, we not only evaluate our method on the B40-C10 on the COCO dataset but also the more challenging B0-C10. As prompt-based methods use a different backbone model compared to our approach, we use the tendency toward an upper bound ($\nabla$) to measure the performance of each method given a specific backbone. Table III illustrates the comparison between our method and prompt-based methods. We observe that our method consistently outperforms prompt-based methods comprehensively, with higher performance across all scenarios, both with and without buffer, compared to the KRT method. Specifically, our approach shows gaps of 7.33% and 4.66% compared to the upper bound, while DualPrompt exhibits gaps of 12.96% and 10.56% compared to the upper bound. In comparison to the Prompt method, we achieve improvements of at least 5.63%. It is worth noting that despite our method having a lower upper bound, the average accuracy on the two benchmarks is improved by at least 4.64%.
Result on VOC benchmark: Tab II summarizes the experimental results on the VOC banchmark. We observe similar trends as seen in the COCO dataset results. Specifically, our method achieves a significant improvement in the final mAP compared to KRT, with an increase of 3.0% (83.4% $\xrightarrow{}$ 86.4%) and 3.0% (80.5% $\xrightarrow{}$ 83.5%), surpassing other rehearsal methods by at least 9.1%. It is noteworthy that KRT experiences a significant drop in accuracy on the more challenging benchmark (B10-C2) without a buffer, while our method maintains good accuracy, resulting in a relative improvement of 3.8%.
@@
TABLE IV: Experimental results (mAP%) of our method and comparison methods on Split-COCO and Split-VOC datasets. All the input size is $448\times 448$, and all metrics are in %. All compared results are obtained from APPLE [12]
Methods Buffer Split-COCO Split-COCO Buffer Split-VOC Split-VOC
Size B40-C10 B0-C20 Size B10-C5 B0-C5
Avg. Acc Last Acc Avg. Acc Last Acc Avg. Acc Last Acc Avg. Acc Last Acc
Upper bound - - 86.43 - 86.43 - - 94.18 - 94.18
FT 0/class 35.83 11.12 ($\downarrow$ 71.92) 51.87 23.60 ($\downarrow$ 58.93) 0/class 74.82 60.09 ($\downarrow$ 32.26) 74.10 59.73 ($\downarrow$ 30.53)
Ours 84.36 82.54 ($\downarrow$ 0.50) 83.39 81.71 ($\downarrow$ 0.82) 94.66 91.98 ($\downarrow$ 0.37) 93.88 89.45 ($\downarrow$ 0.81)
iCaRL [6] 20/class 76.69 65.60 ($\downarrow$ 17.44) 76.53 64.54 ($\downarrow$ 17.41) 5/class 90.78 87.07 ($\downarrow$ 5.28) 88.33 84.78 ($\downarrow$ 5.48)
ER [25] 72.30 64.05 ($\downarrow$ 18.99) 63.59 50.14 ($\downarrow$ 32.81) 86.01 73.91 ($\downarrow$ 18.44) 82.68 68.31 ($\downarrow$ 21.95)
TPCIL [22] 69.41 71.20 ($\downarrow$ 11.84) 73.54 68.89 ($\downarrow$ 14.06) 90.19 84.18 ($\downarrow$ 8.17) 87.95 79.38 ($\downarrow$ 10.88)
PODNet [23] 77.11 66.12 ($\downarrow$ 16.92) 74.78 61.01 ($\downarrow$ 21.94) 90.35 86.35 ($\downarrow$ 6.00) 87.78 84.12 ($\downarrow$ 6.14)
PASS [76] 73.80 59.44 ($\downarrow$ 23.60) 72.16 49.88 ($\downarrow$ 33.07) 86.01 76.93 ($\downarrow$ 15.42) 75.58 51.84 ($\downarrow$ 38.42)
DER++ [24] 66.71 55.77 ($\downarrow$ 27.27) 73.82 67.33 ($\downarrow$ 15.62) 90.22 83.95 ($\downarrow$ 8.40) 88.95 84.82 ($\downarrow$ 5.44)
APPLE [12] 82.05 74.61 ($\downarrow$ 8.43) 83.49 76.65 ($\downarrow$ 6.30) 91.68 89.36 ($\downarrow$ 2.99) 89.52 85.62 ($\downarrow$ 4.64)
Ours 84.79 83.04 ($\downarrow$ 0.00) 83.92 82.53 ($\downarrow$ 0.00) 94.85 92.35 ($\downarrow$ 0.00) 95.62 90.26 ($\downarrow$ 0.00)
Results on Split-COCO and Split-VOC benchmark: Tab IV summarizes the experimental results on the Split-COCO and Split-VOC benchmarks. We observe that our method exhibits significant improvements compared to rehearsal-based methods on these new protocols and achieves higher advancements compared to the SOTA method APPLE in the field of MLCIL. Specifically, our method outperforms APPLE by 8.43% and 6.30% on the two benchmarks in Split-COCO and achieves a maximum improvement of 4.64% on the relatively simpler Split-VOC. Furthermore, even with a buffer size of 0, our method experiences a maximum accuracy drop of only 0.78%, providing further evidence of the effectiveness of our approach.
IV-C Ablation Study
Ablation of Different Components: To demonstrate how each component of the proposed CLIN framework contributes to performance improvements, we conducted ablation studies by building five models on MS-COCO under the B0-C10 benchmark: (1) We introduced a distillation loss as the baseline model. (2) We added a class-independent incremental Network (CINet) to improve the baseline model. (3) We added both the CINet and the proposed token orthogonal loss $\mathcal{L}_{to}$ to improve the baseline model. (4) We added both the CINet and the proposed multi-label contrastive loss $\mathcal{L}_{mc}$ to improve the baseline model. (5) We added all four components to improve the baseline model.
Table V shows the average mAP and each session mAP of the above models. Specifically, the baseline model produced the lowest last mAP value of 58.4%. Adopting the CINet improved the last mAP by 15.0%. Adding the token orthogonal loss $\mathcal{L}_{to}$ further improved the performance by 0.6%. Adding the multi-label contrastive loss $\mathcal{L}_{mc}$ further improved the performance by 1.0%. The final model, which consists of all three components, achieved the best average mAP and last mAP of 79.5% and 74.5%. These experimental results strongly demonstrate that the three components utilized in the proposed CLIN method are highly effective in preventing forgetting and improving performance for MLCIL tasks.
TABLE V: The effectiveness of CLIN’s different components in performance improvements.
Baseline CINet TO loss MC loss Acc (mAP%) in each session Avg.
@@ IV-C Ablation Study
92.6 83.0 79.4 77.3 76.0 75.3 73.8 73.4 ($\uparrow$ 15.0) 78.9
92.8 83.5 79.9 77.4 76.1 75.9 74.3 74.0 ($\uparrow$ 15.6) 79.2
93.0 83.3 79.7 77.5 76.3 76.0 74.5 74.4 ($\uparrow$ 16.0) 79.3
92.9 83.3 79.8 77.9 76.8 76.2 74.6 74.5 ($\uparrow$ 16.1) 79.5
Sensitive Study of Hyper-parameter: The hyper-parameters $\alpha$, $\beta$ and $\lambda$ control the strength of $\mathcal{L}_{mc}$, $\mathcal{L}_{kd}$ and $\mathcal{L}_{to}$ to together control the total optimization objective.

### v2

@@ 4.1 Experimental Setup and Implementation Details
Compared Method. Our method was compared to multiple SOTA continual learning methods following the MLCIL protocol. This included the SOTA MLCIL methods KRT [5], APPLE [33], RebLL [11], and CSC [10], along with various single-label class-incremental learning methods for comparison. These single-label methods comprised regularization-based approaches such as EWC [30], LwF [18], and rehearsal-based methods like ER [28], TPCIL [35], DER++[2], and PODNet[6]. Additionally, we considered prompt-based methods L2P [41], DualPrompt [40], and CODA [31] both utilizing pre-trained ViT-B models and representing the latest SOTA methods in the incremental learning field. The Upper-bound joint trains all classes simultaneously. However, it is not an Incremental Learning (IL) method but serves as an upper bound for performance comparison.
@@
Evaluation metrics. Following the precedent set by prior incremental learning studies [5, 14], we report two key metrics, average accuracy and final accuracy. For a comprehensive evaluation of all learned categories within each session, we utilize the mean average precision (mAP) and report both the average mAP (the mean of the mAP across all sessions) and the final mAP (mAP in the final session). To further enhance the evaluation of performance across all incremental tasks, we include the per-class F1 measure (CF1) and overall F1 measure (OF1) alongside the final accuracy.
@@
Implementation details. All methods were implemented using PyTorch and trained on 2 RTX 3090 GPUs. We employed a TResNetM backbone pre-trained on ImageNet-21k (L2P [41] and Dual-prompt utilized ViT-B/16 pre-trained on ImageNet-21k as the backbone). The input resolution is set to $h\times w=224\times 224$ or $h\times w=448\times 448$ depending on the training protocol. The projection dimension $D$ and hidden dimension $l$ of the CINet module were configured as $768$. Our method was trained for 20 epochs using the Adam optimizer and OneCycleLR scheduler, with a weight decay of $1\times 10^{-4}$. The batch size was set to 64. For training the base model, the learning rate was set to $4\times 10^{-5}$. During the incremental sessions, we set the learning rate to $1\times 10^{-4}$ for COCO and $8\times 10^{-5}$ for VOC. For the hyper-parameters of total optimization, $\alpha$ is typically set to 0.05 and $\beta$ is set to 80. Data augmentation techniques included rand augmentation and cutout. Additionally, we conducted experiments three times and reported the average results.
@@
L2P [41] 73.28 ($\downarrow 6.82$) 67.72 ( $\nabla$ 15.44 ) 73.07 ($\downarrow 7.50$) 70.42 ( $\nabla$ 12.74 )
L2P-R [41] 73.87 ($\downarrow 6.23$) 68.22 ( $\nabla$ 14.94 ) 73.64 ($\downarrow 6.93$) 71.68 ( $\nabla$ 11.48 )
Dual-prompt [40] 74.67 ($\downarrow 5.42$) 69.39 ( $\nabla$ 13.77 ) 74.45 ($\downarrow 6.12$) 71.95 ( $\nabla$ 11.21 )
Dual-prompt-R [40] 74.87 ($\downarrow 5.23$) 70.20 ( $\nabla$ 12.96 ) 74.54 ($\downarrow 6.03$) 72.60 ( $\nabla$ 10.56 )
CODA [31] 75.84($\downarrow 4.26$) 69.87($\nabla$ 5.31) 75.31($\downarrow 5.26$) 72.49($\nabla$ 5.31)
CODA-R [31] 76.07($\downarrow 4.03$) 70.96($\nabla$ 4.22) 75.50($\downarrow 5.07$) 72.92($\nabla$ 4.88)
Ours 78.89 ($\downarrow 1.21$) 72.24 ( $\nabla$ 10.92 ) 79.62 ($\downarrow 0.95$) 76.24 ( $\nabla$ 6.92 )
Ours-R 80.10 ($\downarrow 0.00$) 75.18 ( $\nabla$ 7.98 ) 80.57 ($\downarrow 0.00$) 77.80 ( $\nabla$ 5.36 )
Table 3: Class-Incremental results (mAP%) on MS-COCO dataset against prompt-based CIL method. $\nabla$ indicates the gap towards the Upper Bound of the corresponding backbone. All the input size is $224\times 224$ and all metrics are in %.
   Method     Buffer Size    VOC B0-C4    VOC B10-C2
   Avg. Acc    Last Acc    Avg. Acc    Last Acc
   mAP (%)    mAP (%)    mAP (%)    mAP (%)
   Upper-bound    -    -    93.6    -    93.6
   FT [26]    0    82.4    62.9 ($\downarrow 24.3$)    70.1    43.0 ($\downarrow 38.2$)
   KRT [5]    89.1    80.2 ($\downarrow 7.0$)    84.3    71.5 ($\downarrow 9.7$)
   Ours    91.1    86.2 ($\downarrow 0.0$)    88.7    81.2 ($\downarrow 0.0$)
   iCaRL [25]    2/class    87.2    72.4 ($\downarrow 14.8$)    79.0    66.7 ($\downarrow 16.8$)
   BiC [42]    86.8    72.2 ($\downarrow 15.0$)    81.7    69.7 ($\downarrow 13.8$)
   ER [27]    86.1    71.5 ($\downarrow 15.7$)    81.5    68.6 ($\downarrow 14.9$)
   TPCIL [35]    87.6    77.3 ($\downarrow 9.9$)    80.7    70.8 ($\downarrow 12.7$)
   PODNet [6]    88.1    76.6 ($\downarrow 10.6$)    81.2    71.4 ($\downarrow 12.1$)
   DER++ [2]    87.9    76.1 ($\downarrow 11.1$)    82.3    70.6 ($\downarrow 12.9$)
   KRT [5]    90.7    83.4 ($\downarrow 3.8$)    87.7    80.5 ($\downarrow 3.0$)
   Ours    92.4    87.2 ($\downarrow 0.0$)    90.3    83.5 ($\downarrow 0.0$)
Table 4: Class-incremental results on PASCAL VOC dataset. Compared methods under different protocols against comparison methods. All the input size is $224\times 224$.
4.2 Comparison with State-of-the-arts.
Results on Challenge COCO Protocols: To further examine model robustness under more demanding incremental conditions, we evaluate our method on the B20-C4 and B0-C5 challenge protocols, where the number of incremental stages increases and each session introduces fewer new classes. These conditions exacerbate catastrophic forgetting and intensify both intra-feature and inter-session confusion, making them a more suitable stress test for multi-label incremental learning.
As reported in Table 1, our method achieves the highest performance across all evaluation metrics and significantly surpasses CSC under both challenge configurations. Under B20-C4, the proposed approach achieves a final mAP of 68.4%, exceeding CSC by 7.8%. Under B0-C5, our method reaches a final mAP of 67.9%, improving over CSC by 4.5%. These substantial margins confirm that the CLIN framework is particularly effective in scenarios where old and new labels frequently co-occur within the same image and where incremental stages provide limited opportunities for model adaptation.
The superior performance arises from three core properties of our design: (1) class-level embeddings that eliminate intra-feature mixing, (2) class-independent classifiers that prevent interference between historical and novel concepts, and (3) a contrastive separation mechanism that explicitly mitigates semantic overlap between old and new label spaces. These results strongly validate that the CLIN framework possesses exceptional anti-forgetting ability and sustained learning capability, proving the effectiveness and robustness of our proposed CINet and dual loss mechanism in challenging MLCIL tasks.
@@
Results on Standard COCO Protocols: Table 2 presents the results of the MS-COCO B0-C10 and B40-C10 benchmark. Our method consistently outperforms almost all comparison methods in terms of average accuracy (mAP) and final accuracy (CF1, OF1, and mAP). Specifically, when the buffer size is larger (20 per class), our method achieves the best final accuracies of 75.0% and 77.7% on the two benchmarks, surpassing SOTA CIL methods by 11.9% and 11.4% respectively. It also outperforms the MLCIL method, KRT, by 4.8% and 2.5%. When the buffer size decreases (5 per class), it shows larger performance gains compared to other continuities. Notably, when the buffer size is set to 0, our method maintains superior performance, far exceeding regularization-based methods, and other rehearsal-based methods, with improvements of 42.5% and 45.8% on the B0-C10 benchmark, respectively. Furthermore, our method is more suitable for rehearsal-free scenarios. Specifically, under the B0-C10 protocol, our method achieves average and final accuracy of 78.6% and 71.9%, respectively, compared to KRT (buffer size=0), representing improvements of 4.8% and 6.0%.
Although CSC achieves slightly higher performance than our method on the Standard COCO benchmarks, exceeding it by approximately 0.9% and 2.2%, our approach surpasses CSC by 7.8% and 4.5% under the more demanding Challenge COCO settings. This demonstrates that CLIN exhibits greater stability and superiority in realistic and more challenging incremental learning conditions.
@@
The current SOTA methods in SLCIL are prompt-based approaches. These methods, based on the VIT-B pre-trained model and utilizing prompts, demonstrate powerful capabilities in preventing catastrophic forgetting. For a more comprehensive comparison, we not only evaluate our method on the B40-C10 on the COCO dataset but also the more challenging B0-C10. As prompt-based methods use a different backbone model compared to our approach, we use the tendency towards an upper bound ($\nabla$) to measure the performance of each method given a specific backbone. Table 3 illustrates the comparison between our method and prompt-based methods. We observe that our method consistently outperforms prompt-based methods comprehensively, with higher performance across all scenarios, both with and without a buffer, compared to the KRT method. Specifically, our approach shows gaps of 7.33% and 4.66% compared to the upper bound, while DualPrompt exhibits gaps of 12.96% and 10.56% compared to the upper bound. In comparison to the Prompt method, we achieve improvements of at least 5.63%. It is worth noting that despite our method having a lower upper bound, the average accuracy on the two benchmarks is improved by at least 4.64%.
   Methods    Buffer    Split-COCO    Split-COCO
   Size    B40-C10    B0-C20
   Avg. Acc    Last Acc    Avg. Acc    Last Acc
   Upper bound    -    -    86.43    -    86.43
   FT    0/class    35.83    11.12 ($\downarrow$ 71.92)    51.87    23.60 ($\downarrow$ 58.93)
   Ours    84.36    82.54 ($\downarrow$ 0.50)    83.39    81.71 ($\downarrow$ 0.82)
   iCaRL [25]    20/class    76.69    65.60 ($\downarrow$ 17.44)    76.53    64.54 ($\downarrow$ 17.41)
   ER [27]    72.30    64.05 ($\downarrow$ 18.99)    63.59    50.14 ($\downarrow$ 32.81)
   TPCIL [35]    69.41    71.20 ($\downarrow$ 11.84)    73.54    68.89 ($\downarrow$ 14.06)
   PODNet [6]    77.11    66.12 ($\downarrow$ 16.92)    74.78    61.01 ($\downarrow$ 21.94)
   PASS [49]    73.80    59.44 ($\downarrow$ 23.60)    72.16    49.88 ($\downarrow$ 33.07)
   DER++ [2]    66.71    55.77 ($\downarrow$ 27.27)    73.82    67.33 ($\downarrow$ 15.62)
   APPLE [33]    82.05    74.61 ($\downarrow$ 8.43)    83.49    76.65 ($\downarrow$ 6.30)
   Ours    84.79    83.04 ($\downarrow$ 0.00)    83.92    82.95 ($\downarrow$ 0.00)
@@
Table 5: Experimental results (mAP%) of our method and comparison methods on Split-COCO datasets. All the input size is $448\times 448$ and all metrics are in %. All compared results are obtained from APPLE [33]
   Methods    Buffer    Split-VOC    Split-VOC
   Size    B10-C5    B0-C5
   Avg. Acc    Last Acc    Avg. Acc    Last Acc
   Upper bound    -    -    94.18    -    94.18
   FT    0/class    74.82    60.09 ($\downarrow$ 32.26)    74.10    59.73 ($\downarrow$ 30.53)
   Ours    94.66    91.98 ($\downarrow$ 0.37)    93.88    89.45 ($\downarrow$ 0.81)
   iCaRL [25]    5/class    90.78    87.07 ($\downarrow$ 5.28)    88.33    84.78 ($\downarrow$ 5.48)
   ER [27]    86.01    73.91 ($\downarrow$ 18.44)    82.68    68.31 ($\downarrow$ 21.95)
   TPCIL [35]    90.19    84.18 ($\downarrow$ 8.17)    87.95    79.38 ($\downarrow$ 10.88)
   PODNet [6]    90.35    86.35 ($\downarrow$ 6.00)    87.78    84.12 ($\downarrow$ 6.14)
   PASS [49]    86.01    76.93 ($\downarrow$ 15.42)    75.58    51.84 ($\downarrow$ 38.42)
   DER++ [2]    90.22    83.95 ($\downarrow$ 8.40)    88.95    84.82 ($\downarrow$ 5.44)
   APPLE [33]    91.68    89.36 ($\downarrow$ 2.99)    89.52    85.62 ($\downarrow$ 4.64)
   Ours    94.85    92.35 ($\downarrow$ 0.00)    95.62    90.26 ($\downarrow$ 0.00)
Table 6: Experimental results (mAP%) of our method and comparison methods on Split-VOC datasets. All the input size is $448\times 448$ and all metrics are in %. All compared results are obtained from APPLE [33]
Baseline CINet MC loss Acc (mAP%) in each session Avg.
@@
92.6 83.0 79.4 77.7 76.4 76.1 74.6 73.7 ($\uparrow$ 15.3) 79.2
93.0 83.4 79.9 78.3 77.0 76.7 75.5 75.0 ($\uparrow$ 16.6) 79.8
Table 7: The effectiveness of CLIN’s different components in performance improvements.
[…]
Results on Split-COCO and Split-VOC benchmark: Table 5 and Table 6 summarize the experimental results on the Split-COCO and Split-VOC benchmarks. We observe that our method exhibits significant improvements compared to rehearsal-based methods on these new protocols and achieves higher advancements compared to the SOTA method APPLE in the field of MLCIL. Specifically, our method outperforms APPLE by 8.43% and 6.30% on the two benchmarks in Split-COCO and achieves a maximum improvement of 4.64% on the relatively simpler Split-VOC. Furthermore, even with a buffer size of 0, our method experiences a maximum accuracy drop of only 0.78%, providing further evidence of the effectiveness of our approach.
We observe that the CLIN method achieves significant performance gains on the MS-COCO benchmark, yet the relative margin of improvement is less pronounced on PASCAL VOC. We attribute this to the differential suitability between the core mechanism of our method and the specific characteristics of the two datasets. COCO is characterized by high density and complex scenes, which lead to severe Intra-Feature Confusion—the very problem where CLIN’s class-level decoupling strategy is designed to be maximally effective. In contrast, VOC scenes are simpler and feature lower object density, resulting in less intrinsic feature entanglement. In these low-confusion scenarios, the advantage of our fine-grained, Token-based decoupling is diminished, suggesting that CLIN is optimally suited for tackling high-feature-confusion problems in complex multi-label environments.
4.3 Ablation Study
Ablation of Different Components: To demonstrate how each component of the proposed CLIN framework contributes to performance improvements, we conducted ablation studies by building four models on MS-COCO under the B0-C10 benchmark: (1) We introduced a distillation loss as the baseline model. (2) We added a class-independent incremental Network (CINet) to improve the baseline model. (3) We added both the CINet and the proposed multi-label contrastive loss $\mathcal{L}_{mc}$ to improve the baseline model. (4) We added all three components to improve the baseline model.
[…]